# SynthSoM: A synthetic intelligent multi-modal sensing-communication dataset for Synesthesia of Machines (SoM)

### v1

@@ Abstract
  • The SynthSoM dataset encompasses multiple data modalities, including RF communications, i.e., 373K sets of channel matrices and 104K sets of path loss, RF sensing, i.e., 230K sets of mmWave radar waveforms with 61K radar point clouds, and non-RF sensing, i.e., 455K RGB images, 891K depth maps, and 241K sets of LiDAR point clouds.
@@ Methods
To generate and collect the non-RF sensory data, i.e., RGB images, depth maps, and LiDAR point clouds, we utilize an open-sourced software developed by Microsoft, i.e., AirSim [22]. AirSim provides a virtual 3D environment to simulate diverse scenarios and conditions. First, by exploiting a city modeling software, i.e., CityEngine [23], we utilize satellite maps in the real world to generate building models, which are exported to Unreal Engine 4 (UE4) [24]. A powerful software, i.e., RoadRunner [25], is further utilized to conduct detailed processing on road layout, terrain features, and plants, which are also exported to UE4. Furthermore, we place intelligent agents, such as UAVs and vehicles (buses and cars), within the scenario. Finally, we refine the scenario details in UE4 and perform in-depth rendering. In the constructed scenario, the intelligent agent and roadside facility (RSF) are equipped with multi-modal sensors, including RGB cameras, depth cameras, and LiDAR devices. According to the requirement of different tasks, key parameters of multi-modal sensors, such as field-of-view (FoV), resolution, LiDAR range, and the number of LiDAR channels, can be adjusted.
@@
To generate and collect the RF sensory data, i.e., mmWave radar information, we utilize efficient commercial software developed by REMCOM, i.e., WaveFarer [26]. WaveFarer exploits the advanced computational approach, such as full-wave electromagnetic simulation, to characterize wave propagation. First, we import 3D scenario models from the project of AirSim into WaveFarer to achieve the in-depth integration between AirSim and WaveFarer. Second, we set materials for 3D scenario models and configure the electrical conductivity and magnetic permeability of materials. Third, we configure the diffuse scattering parameter to ensure that the mmWave radar can detect objects. Finally, we import intelligent agent models into WaveFarer and set their electromagnetic materials. In the constructed scenario, intelligent agents and RSF are equipped with mmWave radars, which leverage the linear frequency modulated continuous wave (FMCW) operating in the frequency range of $77$ GHz to $81$ GHz. Key parameters of mmWave radars, e.g., detection range, angular resolution, and sweep frequency range of FMCW, can be adjusted.
@@
To generate and collect communication information, we utilize efficient commercial software developed by REMCOM, i.e., Wireless InSite [27]. Based on the powerful ray-tracing and diffraction algorithm, Wireless InSite can imitate realistic wireless propagation effects. First, we import models in AirSim into Wireless InSite, thus achieving the in-depth integration between AirSim and Wireless InSite. Then, we set materials of different types of models. Moreover, we conduct the double size operation on imported models to enhance the completeness of the triangle mesh. Finally, we place the intelligent agent within the scenario and set the frequency band, antenna type, and waveform type. In the constructed scenario, intelligent agents and RSF are equipped with communication equipment. Key parameters of communication equipment, such as carrier frequency, bandwidth, waveform type, and antenna type, can also be adjusted.
@@
To achieve the precise alignment, the high-fidelity scenarios in AirSim, WaveFarer, and Wireless InSite are accurately aligned. Specifically, the size/coordinate of each object model in each software at the initial snapshot is the same. Currently, the SynthSoM dataset covers air-ground multi-link cooperative scenarios with five typical areas, i.e., urban crossroad, urban wide lane, urban overpass, suburban fork, and mountain road. For clarity, taking AirSim as an example, Figure 2 shows the existing air-ground multi-link cooperative scenarios in the SynthSoM dataset.
@@
Figure 2: Air-ground multi-link cooperative scenarios in the SynthSoM dataset.
@@
In AirSim, Python is utilized to control the amount of rain and snow for the weather simulation. The amount of rain, degree of water accumulation, amount of snow, and degree of snow accumulation are represented by Rain, Roadwetness, Snow, and RoadSnow, respectively. By setting Rain, Roadwetness, Snow, and RoadSnow, sunny, rainy, and snowy days can be mimicked. Nonetheless, because of the inaccurate characterization of object properties in AirSim, e.g., material reflectivity, weather conditions solely affect RGB and depth cameras and not LiDAR devices. This limitation is also present in other software, e.g., CARLA [28] and LGSVL [29]. To overcome this drawback, we exploit a LiDAR point weather augmentation method, i.e., LISA [30]. The LISA method develops a physics-based hybrid Monte Carlo based LiDAR scatterer simulator for complex weather conditions, e.g., rainy and snowy days. The rain rate and the snow rate in the LISA method are set to $50$ mm/hr and $10$ mm/hr, respectively [30]. In addition to different weather conditions, different times of the day are also taken into account. By adjusting the brightness of the sky lighting and the position of the sun, we mimic different times of the day, such as morning and night. Finally, by placing different numbers of intelligent agents, we simulate scenarios with high, medium, and low intelligent agent densities.
@@
In WaveFarer, for the weather simulation, we utilize a single raindrop/snowflake model and import it into WaveFarer to check if its sizes match the current scenario. By further leveraging an open-sourced 3D computer graphics software, i.e., Blender [31], we scale the individual raindrop/snowflake model to an appropriate size and combine it to create a rain/snow model. Next, we import the processed rain and snow models into WaveFarer to simulate rainy and snowy days, respectively. Then, the temperature and humidity are set. For the rainy day, the temperature is set to $22.2$ degrees Celsius and the humidity is set to $100$%. For the snowy day, the temperature is set to $-10$ degrees Celsius and the humidity is set to $20$%. To mimic high, medium, and low intelligent agent densities, we place different numbers of intelligent agents. For the setting of the antenna type, we design an antenna type and set the $3$ dB attenuation and cutoff beamwidth in the vertical and horizontal directions. Transceivers are configured based on the desired array type and the transmit power of mmWave radar is set. In the FMCW setting, parameters such as the start and stop frequencies of the scan, the duration of a single scan, and the number of scans per snapshot are selected.
@@
In Wireless InSite, we import the rain and snow models that are constructed in WaveFarer into Wireless InSite to mimic rainy and snowy days, respectively. The temperature and humidity for rainy and snowy days in Wireless InSite are the same as those for rainy and snowy days in WaveFarer. Since the intelligent agent density has a significant impact on the channel data [32, 33], we also consider high, medium, and low intelligent agent densities by placing different numbers of intelligent agents. Furthermore, in the Waveforms module of Wireless InSite, we can select the different waveform types and set different frequency bands, such as sub-6 GHz and mmWave. In the Antennas module of Wireless InSite, we can configure different antenna types, including single-input single-output (SISO), multiple-input multiple-output (MIMO), and massive MIMO.
@@
To achieve the precise alignment, the weather conditions and intelligent agent densities in AirSim, WaveFarer, and Wireless InSite are aligned by properly setting the weather parameter and the number of intelligent agents, respectively. As a result, the multi-modal sensory data and communication data are aligned under different scenario conditions. Currently, the SynthSoM dataset covers three weather conditions, i.e., sunny, rainy, and snowy days, two times of the day, i.e., morning and night, three types of intelligent agent densities, i.e., high, medium, and low, two frequency bands, i.e., sub-6 GHz and mmWave, and three antenna types, i.e., SISO, MIMO, and massive MIMO. For clarity, taking the urban wide lane scenario as an example, Figure 3 illustrates comprehensive scenario conditions in the SynthSoM dataset. Note that the simultaneous incorporation and interaction of various scenario conditions enhance the comprehensiveness of the SynthSoM dataset.
@@
Figure 3: Comprehensive conditions in the SynthSoM dataset taking the urban wide lane scenario as an example.
@@
To generate the realistic trajectory of the dynamic intelligent agent, we exploit an open-sourced microscopic traffic software, i.e., Simulation of Urban MObility (SUMO) [34]. First, we export the road network from RoadRunner and import it into SUMO to obtain the initial geometric structure of the road network, which is further revised according to road conditions. Second, we select several positions as anchor points in the road network of RoadRunner and SUMO and measure the coordinates of these anchor points. By calculating the average of the affine transformation matrices and performing translation, rotation, and scaling, we align the coordinate systems of RoadRunner and SUMO. Third, we set traffic rules for each lane, such as straight, left turn, and right turn. Meanwhile, we install traffic lights at each intersection and differentiate lanes with solid and dashed lines. Fourth, according to scenario requirements, we calculate simulation parameters, generate random traffic flows, and correct abnormal traffic flows. As a result, the trajectory of the dynamic intelligent agent is generated in SUMO. Fifth, we import the trajectory generated in SUMO into UE4. Similar to the coordinate alignment operation between RoadRunner and SUMO, we also achieve coordinate alignment between UE4 and SUMO. Finally, trajectories generated in UE4 are imported into AirSim, Wireless InSite, and WaveFarer to ensure that the three software have consistent trajectories.
@@
In AirSim and WaveFarer, as previously mentioned, trajectories of dynamic intelligent agents are set. Through the predefined trajectory, multi-modal sensors mounted on dynamic intelligent agents collect and automatically export multi-modal sensory data, including RGB images, depth maps, LiDAR point clouds, and mmWave radar point clouds. In Wireless InSite, the generated communication scenarios are queued for simulation by a script. For the SISO condition, channel data, such as path loss and channel matrix, can be directly collected and exported. However, for the MIMO or the massive MIMO condition, Wireless InSite cannot directly export the collected channel data and instead requires manual exportation. To overcome this limitation, we utilize Python to automatically control the mouse to export the collected channel data without manual exportation. The SynthSoM dataset encompasses multiple data modalities, including RF channel large-scale and small-scale fading data, RF mmWave radar sensory data, and non-RF sensory data, e.g., RGB images, depth maps, and LiDAR point clouds.
In summary, the developed platform integrates three high-precision software, i.e., AirSim, WaveFarer, and Wireless InSite, and further achieves in-depth integration and precise alignment of them. There are four steps involved in developing the simulation platform, including high-fidelity scenario construction, comprehensive scenario condition simulation, dynamic scenario generation, and data collection and export. Based on the developed platform, the SynthSoM dataset is constructed, which encompasses multiple data modalities under various air-ground multi-link cooperative scenarios with comprehensive conditions. For clarity, more detailed information on the SynthSoM dataset, such as the installation of software, the description of each scenario, the trajectory of intelligent agent, etc., is available on GitHub (https://github.com/ZiweiHuang96/SynthSoM).
@@
To utilize the SynthSoM dataset for communication and multi-modal sensing tasks, we divide the SynthSoM dataset into five folders according to different scenarios. The scenario folder contains different conditions, and further contains different data modalities, i.e., RF sensing, non-RF sensing, and RF communications. The data directory hierarchy and content of the SynthSoM dataset are shown in Fig. 5. Due to the significantly massive data amount of the SynthSoM dataset, i.e., approximately $3$ TB, we select a specific scenario condition, i.e., data collected by vehicles in the urban wide lane scenario with low intelligent agent density on a sunny morning, to upload to Figshare in zip compression format. The remaining scenario conditions in the SynthSoM dataset are fully open-sourced and uploaded to OneDrive in zip compression format, as listed in Table 2.
Scenario Download Link
Air-Ground Multi-Link https://1drv.ms/f/c/2ecb56630cb1c9b9/Eqa557lsp3JGkec3RReQmfgBLXnro3salsWjtZcP3-HbPA?e=rxHe4K
Cooperative Urban Crossroad
Air-Ground Multi-Link https://1drv.ms/f/c/0dac85eca919bb25/EruKFMFBgkhDiKWK_8BLByQBVOoshS0gnrgzXwBTVnxt_Q?e=Dlyuga
Cooperative Urban Wide Lane
Air-Ground Multi-Link https://1drv.ms/f/c/e437ff297b90b20a/EmHAx7MtufNIgCnfI5sBA_0BNxZaH8A1aLqf2y4rIYpfsQ?e=k83QrC
Cooperative Suburban Fork
Air-Ground Multi-Link https://1drv.ms/f/c/49c410559939c091/Ek34R3zFUDpGlMeUyyKupXcBSFdcU0DpNEypzHSDNCdpAg?e=efCB2R
Cooperative Urban Overpass
Air-Ground Multi-Link https://1drv.ms/f/c/99f7731061a2474a/ElYoKcTXYtZNm5O6z9J91m0BWkuXNKngm6-5KC2fvLIx6A?e=JsFInZ
Cooperative Mountain Road
Table 2: SynthSoM dataset in OneDrive.
@@
Overall, the SynthSoM dataset contains 373K sets of channel matrices, 104K sets of path loss, 230K sets of mmWave radar waveforms with 61K radar point clouds, 455K RGB images, 891K depth maps, and 241K sets of LiDAR point clouds. The open-sourced SynthSoM dataset can provide consistent data for SoM-related algorithm cross comparison, model calibration, and baseline implementation.
@@
Figure 6: Visualization and statistical properties of RF channels under different conditions. (a) Path loss heatmap in urban crossroad scenario with sub-6 GHz frequency band. (b) Path loss heatmap in urban crossroad scenario with mmWave frequency band. (c) Propagation paths in mountain road scenario under SISO. (d) Propagation paths in mountain road scenario under MIMO. (e) Propagation paths in mountain road scenario under massive MIMO. (f) SCCFs in mountain road scenario with sub-6 GHz and mmWave frequency bands under MIMO and massive MIMO conditions.

### v2

@@ Abstract
  • The SynthSoM dataset encompasses multiple data modalities, including RF communications, i.e., 140K sets of channel matrices and 18K sets of path loss, RF sensing, i.e., 136K sets of mmWave radar waveforms with 38K radar point clouds, and non-RF sensing, i.e., 145K RGB images, 290K depth maps, and 79K sets of LiDAR point clouds.
@@ Methods
To generate and collect the non-RF sensory data, i.e., RGB images, depth maps, and LiDAR point clouds, we utilize an open-sourced software developed by Microsoft, i.e., AirSim [22]. AirSim provides a virtual 3D environment to simulate diverse scenarios and conditions. First, by exploiting a city modeling software, i.e., CityEngine [23], we utilize satellite maps in the real world to generate building models, which are exported to Unreal Engine 4 (UE4) [24]. A powerful software, i.e., RoadRunner [25], is further utilized to conduct detailed processing on road layout, terrain features, and plants, which are also exported to UE4. Furthermore, we place intelligent agents, such as UAVs and vehicles (buses and cars), within the scenario. Finally, we refine the scenario details in UE4 and perform in-depth rendering. In the constructed scenario, the intelligent agent and roadside facility (RSF) are equipped with multi-modal sensors, including RGB cameras, depth cameras, and LiDAR devices. First, to collect RGB images, AirSim utilizes the rendering pipeline of UE4 to simulate camera imaging. Each simulated camera is equipped with a SceneCaptureComponent2D component from UE4, which captures the scenario from the camera frustum and renders the view. Second, to collect depth maps, UE4 generates a depth buffer of the scenario for each snapshot during rendering. Through the SceneCapture component, the depth value corresponding to each pixel within the camera frustum can be captured. Third, to collect LiDAR point clouds, AirSim simulates a LiDAR by emitting a multitude of laser rays into space and detecting the intersection points of these rays with objects to obtain distance points. In addition, AirSim captures the occlusion, reflections, and sensor noise/failure by utilizing Z-Buffer method to achieve pixel-level occlusion effects, exploiting physically-based reflection properties, and adding noise models to sensors to simulate real-world measurement inaccuracies, respectively. According to the requirement of different tasks, key parameters of multi-modal sensors, such as field-of-view (FoV), resolution, LiDAR range, and the number of LiDAR channels, can be adjusted.
@@
To generate and collect the RF sensory data, i.e., mmWave radar information, we utilize efficient commercial software developed by REMCOM, i.e., WaveFarer [26]. WaveFarer exploits the advanced computational approach, such as full-wave electromagnetic simulation, to characterize wave propagation. First, we import 3D scenario models from the project of AirSim into WaveFarer to achieve the in-depth integration between AirSim and WaveFarer. Specifically, we establish the units for 3D scenario model importation to ensure the accuracy of the scenario dimensions. Subsequently, we import the STereoLithography (STL) format models provided by AirSim and assign the corresponding material properties, such as wood, metal, and concrete, according to the object categories, such as trees, vehicles, and buildings. Second, we set materials for 3D scenario models and configure the electrical conductivity and magnetic permeability of materials. Third, we configure the diffuse scattering parameter to ensure that the mmWave radar can detect objects. Finally, we import intelligent agent models into WaveFarer and set their electromagnetic materials. In the constructed scenario, intelligent agents and RSF are equipped with mmWave radars, which leverage the linear frequency modulated continuous wave (FMCW) operating in the frequency range of $77$ GHz to $81$ GHz. The sampling interval for mmWave radar information is $5$ nanoseconds, with each chirp containing $1000$ sampling points and each snapshot consisting of $101$ chirps, resulting in a waveform file length of $101,000$ samples per snapshot. Furthermore, to collect mmWave radar information, the “Create Linear Chirp Simulation” script provided by WaveFarer is utilized to perform ray-tracing simulation for FMCW radar and to acquire the echo signals. Then, the official “Generate Range-Doppler” script is employed to process the echo signals obtained from the previous step, simulating the down-conversion and filtering operations of an actual FMCW radar. Next, a three-dimensional fast Fourier transform (3D-FFT) is exploited to the processed echo signals across the range, velocity, and angle dimensions, followed by the implementation of constant false alarm rate (CFAR) detection to obtain the mmWave radar point cloud data. Additionally, WaveFarer captures the occlusion, reflections, and sensor noise/failure by importing 3D models, adjusting electromagnetic parameters of object materials, and adjusting the detection threshold of the CFAR algorithm to characterize different noise levels at the code level, respectively. Key parameters of mmWave radars, e.g., detection range, angular resolution, and sweep frequency range of FMCW, can be adjusted.
@@
To generate and collect communication information, we utilize efficient commercial software developed by REMCOM, i.e., Wireless InSite [27]. Based on the powerful ray-tracing technology, Wireless InSite can imitate realistic wireless propagation effects, where RF interference and multipath propagation effects can be captured. For the capturing of RF interference, the ray-tracing technology simulates propagation paths of electromagnetic waves by integrating reflection, refraction, diffraction, and scattering models. The utilization of ray-tracing technology can facilitate comprehensively analyzing the source of RF interference and examining the interactions between different signal sources, thus evaluating the effect of RF interference. For the capturing of multipath propagation effects, a set of rays can be generated from the transmitter (Tx) with a specific angular separation. Then, rays are propagated in the environment up to a maximum number of interactions with surfaces, and are considered to reach the receiver (Rx) if rays meet receiving conditions. First, we import models in AirSim into Wireless InSite to achieve the in-depth integration between AirSim and Wireless InSite. Similar to the previously mentioned in-depth integration with AirSim and WaveFarer, the STL format 3D scenario models provided by AirSim are imported and the corresponding material settings, e.g., wood, metal, and concrete, are applied according to the object categories, e.g., trees, vehicles, buildings. Then, we set materials of different types of models. Moreover, we conduct the double size operation on imported models to enhance the completeness of the triangle mesh. Finally, we place the intelligent agent within the scenario and set the frequency band, antenna type, and waveform type. In the constructed scenario, intelligent agents and RSF are equipped with communication equipment. Key parameters of communication equipment, such as carrier frequency, bandwidth, waveform type, and antenna type, can also be adjusted. By utilizing the ray-tracing technology in different environments with different 3D scenario models, the RF interference and multipath propagation effects across different environments can be properly captured. Also, Wireless InSite captures the occlusion, reflections, and sensor noise/failure by simulating the blocking of signals via obstacles, utilizing shooting and bouncing ray and the uniform theory of diffraction methods, and providing hardware noise modeling functions, respectively.
Note that a detailed parameter setting in the first step, i.e., high-fidelity scenario construction, can be observed in https://github.com/ZiweiHuang96/SynthSoM. For example, for Scenario 1, i.e., Air-Ground Multi-Link Cooperative Urban Crossroad, key parameters, including sensor position, sensor parameter, and RSF position, can be found in https://github.com/ZiweiHuang96/SynthSoM/blob/main/urban_crossroad/README.md.
@@
To achieve the precise alignment, the high-fidelity scenarios in AirSim, WaveFarer, and Wireless InSite are accurately aligned. The size/coordinate of each object model in each software at the initial snapshot is the same. Specifically, when exporting scenario models from UE4 to WaveFarer and Wireless InSite, the STL format is utilized for export, ensuring no rotation, scaling, or origin offset. By further aligning the coordinate systems, verifying a $1:1$ scale, and unifying the world origin, the strict consistency with the scenario models in UE4 in terms of geometric coordinates, scale, and position is guaranteed. Currently, the SynthSoM dataset covers air-ground multi-link cooperative scenarios with five typical areas, i.e., urban crossroad, urban wide lane, urban overpass, suburban fork, and mountain road. For clarity, taking AirSim as an example, Figure 2 shows the existing air-ground multi-link cooperative scenarios in the SynthSoM dataset.
@@
Figure 2: Air-ground multi-link cooperative scenarios in the SynthSoM dataset. (a) Urban crossroad scenario. (b) Urban wide lane scenario. (c) Urban overpass scenario. (d) Suburban fork scenario. (e) Mountain road scenario.
@@
In AirSim, Python is utilized to control the amount of rain and snow for the weather simulation. The amount of rain, degree of water accumulation, amount of snow, and degree of snow accumulation are represented by Rain, Roadwetness, Snow, and RoadSnow, respectively. By setting Rain, Roadwetness, Snow, and RoadSnow, sunny, rainy, and snowy days can be mimicked. To simulate rainy and snowy conditions on RGB images and depth maps, the RGB images and depth maps are captured by the AirSim plugin based on the actual scenario weather, meaning that the RGB image is a snapshot of the scenario instantaneous state, and the depth map is a snapshot of the scenario instantaneous state. Nonetheless, because of the inaccurate characterization of object properties in AirSim, e.g., material reflectivity, weather conditions solely affect RGB and depth cameras and not LiDAR devices. This limitation is also present in other software, e.g., CARLA [28] and LGSVL [29]. To overcome this drawback, we exploit a LiDAR point weather augmentation method, i.e., LISA [30]. The LISA method simulates the propagation process of laser beams in the physical world to determine the reflection positions and corresponding power intensities. Specifically, the Monte Carlo algorithm is utilized to simulate the distribution of scatterers, such as rain and snow particles in space, the Fresnel reflection is employed to depict the propagation path of light in space, and the position and intensity of the scatterers are measured based on the minimum detectable power of the sensor. The rain rate and the snow rate in the LISA method are set to $50$ mm/hr and $10$ mm/hr, respectively [30]. In addition to different weather conditions, AirSim is also capable of simulating various times of the day by adjusting the sun altitude and the intensity of skylight. The angle of the sun altitude influences the angle of light incidence, thus capturing the distinct effects of dawn, morning, and night. Furthermore, the intensity of skylight directly reflects the global environmental illumination and can be matched with different values to represent the varying effects of dawn, morning, and night. For the simulation of dawn, i.e., dim light during the day, the sun elevation angle $Sun_{\mathrm{height}}$ is set to 00 with the sun axis aligned parallel to the horizon $Elevation$ also set to 00. The sky light parameter $Sky_{\mathrm{light}}$ is set to $1.0$, which means base ambient illumination. For the simulation of morning, i.e., strong light during the day, the sun elevation angle $Sun_{\mathrm{height}}$ is set to $1.0$ with the sun axis aligned parallel to the horizon $Elevation$ set to $90^{\circ}$. The sky light parameter $Sky_{\mathrm{light}}$ is set to $2.0$, which means enhanced global illumination. For the simulation of night, i.e., nighttime environment, the sun elevation angle $Sun_{\mathrm{height}}$ is set to $-1$. According to the astronomical twilight criteria, the sun axis aligned parallel to the horizon $Elevation$ is set to $-18^{\circ}$, where the sun is below the horizon. The sky light parameter $Sky_{\mathrm{light}}$ is set to $0.2$, which is the minimum ambient light. Furthermore, the starlight brightness parameter $Star_{\mathrm{intensity}}$ is set to $0.8$. Finally, by placing different numbers of intelligent agents, we simulate scenarios with high, medium, and low intelligent agent densities.
@@
To achieve the in-depth integration between AirSim and WaveFarer, we utilize a single rain/snow model and import it into WaveFarer to check if its sizes match the current scenario. By further leveraging an open-sourced 3D computer graphics software, i.e., Blender[31], we scale the individual rain/snow model to an appropriate size and combine it to create a rain/snow model. Next, we import the processed rain and snow models into WaveFarer to simulate rainy and snowy days, respectively. Then, the temperature and humidity are set. For the rainy day, the temperature is set to $22.2$ degrees Celsius and the humidity is set to $100$%. For the snowy day, the temperature is set to $-10$ degrees Celsius and the humidity is set to $20$%. To mimic high, medium, and low intelligent agent densities, we place different numbers of intelligent agents. For the setting of the antenna type, we design an antenna type and set the $3$ dB attenuation and cutoff beamwidth in the vertical and horizontal directions. Transceivers are configured based on the desired array type and the transmit power of mmWave radar is set. In the FMCW setting, parameters such as the start and stop frequencies of the scan, the duration of a single scan, and the number of scans per snapshot are selected.
@@
To achieve the in-depth integration between AirSim and Wireless InSite, we import the rain and snow models that are constructed in WaveFarer into Wireless InSite to mimic rainy and snowy days, respectively. The temperature and humidity for rainy and snowy days in Wireless InSite are the same as those for rainy and snowy days in WaveFarer. Since the intelligent agent density has a significant impact on the channel data [32, 33], we also consider high, medium, and low intelligent agent densities by placing different numbers of intelligent agents. Furthermore, in the Waveforms module of Wireless InSite, we can select the different waveform types and set different frequency bands, such as sub-6 GHz and mmWave. In the Antennas module of Wireless InSite, we can configure different antenna types, including single-input single-output (SISO), multiple-input multiple-output (MIMO), and massive MIMO.
@@
To achieve the precise alignment of AirSim, WaveFarer, and Wireless InSite, the weather conditions and intelligent agent densities in the aforementioned software are aligned by properly setting the weather parameter and the number of intelligent agents, respectively. As a result, the multi-modal sensory data and communication data are aligned under different scenario conditions. Currently, the SynthSoM dataset covers three weather conditions, i.e., sunny, rainy, and snowy days, two times of the day, i.e., morning and night, three types of intelligent agent densities, i.e., high, medium, and low, two frequency bands, i.e., sub-6 GHz and mmWave, and three antenna types, i.e., SISO, MIMO, and massive MIMO. For clarity, taking the urban wide lane scenario as an example, Figure 3 illustrates comprehensive scenario conditions in the SynthSoM dataset. Note that the simultaneous incorporation and interaction of various scenario conditions enhance the comprehensiveness of the SynthSoM dataset.
@@
Figure 3: Comprehensive conditions in the SynthSoM dataset taking the urban wide lane scenario as an example. (a) Sunny day. (b) Rainy day. (c) Snowy day. (d) Morning. (e) Night. (f) High intelligent agent density. (g) Medium intelligent agent density. (h) Low intelligent agent density. (i) Sub-6 GHz band. (j) mmWave band. (k) SISO condition. (l) MIMO condition. (m) Massive MIMO condition.
@@
To generate the realistic trajectory of the dynamic intelligent agent, we exploit an open-sourced microscopic traffic software, i.e., Simulation of Urban MObility (SUMO) [34]. First, we export the road network from RoadRunner and import it into SUMO to obtain the initial geometric structure of the road network, which is further revised according to road conditions. Second, we select several positions as anchor points in the road network of RoadRunner and SUMO and measure the coordinates of these anchor points. By calculating the average of the affine transformation matrices and performing translation, rotation, and scaling, we align the coordinate systems of RoadRunner and SUMO. Third, we set traffic rules for each lane, such as straight, left turn, and right turn. Meanwhile, we install traffic lights at each intersection and differentiate lanes with solid and dashed lines. Fourth, according to scenario requirements, we calculate simulation parameters, generate random traffic flows, and correct abnormal traffic flows. As a result, the trajectory of the dynamic intelligent agent is generated in SUMO. Fifth, we import the trajectory generated in SUMO into UE4, which can be further imported into AirSim, WaveFarer, and Wireless InSite, thus achieving an in-depth integration among them. Similar to the coordinate alignment operation between RoadRunner and SUMO, we also achieve coordinate alignment between UE4 and SUMO. Finally, trajectories generated in UE4 are imported into AirSim, Wireless InSite, and WaveFarer to ensure that the three software have consistent trajectories, thus achieving a precise alignment among them. Note that a detailed parameter setting in the third step, i.e., dynamic scenario generation, can be observed in https://github.com/ZiweiHuang96/SynthSoM. For example, for Scenario 1, i.e., Air-Ground Multi-Link Cooperative Urban Crossroad, the parameter related to intelligent agent trajectory is given and can be found in https://github.com/ZiweiHuang96/SynthSoM/blob/main/urban_crossroad/README.md. Therefore, for the data annotation, the label or ground truth of position and orientation parameters related to intelligent agents and RSFs is naturally contained in the synthetic SynthSoM dataset. For vehicles, they accelerate, decelerate, and change lanes on a two-dimensional (2D) ground, thus lacking height information. Different from vehicles, UAVs fly in 3D space, possessing altitude information and exhibiting ascending and descending behaviors. Furthermore, different scenarios also present entirely different trajectories for vehicles and UAVs. Therefore, vehicles and UAVs have different trajectories, and the trajectories vary significantly across different scenarios, thus enhancing the diversity of the SynthSoM dataset.
@@
In AirSim and WaveFarer, as previously mentioned, trajectories of dynamic intelligent agents are set. Through the predefined trajectory, multi-modal sensors mounted on dynamic intelligent agents collect and automatically export multi-modal sensory data, including RGB images, depth maps, LiDAR point clouds, and mmWave radar point clouds. In Wireless InSite, the generated communication scenarios are queued for simulation by a script. For the SISO condition, channel data, such as path loss and channel matrix, can be directly collected and exported. However, for the MIMO or the massive MIMO condition, Wireless InSite cannot directly export the collected channel data and instead requires manual exportation. To overcome this limitation, we utilize Python to automatically control the mouse to export the collected channel data without manual exportation. The SynthSoM dataset encompasses multiple data modalities, including RF channel large-scale and small-scale fading data, RF mmWave radar sensory data, and non-RF sensory data, e.g., RGB images, depth maps, and LiDAR point clouds. Due to the in-depth integration and precise alignment of the aforementioned three steps, i.e., high-fidelity scenario construction, comprehensive scenario condition simulation, and dynamic scenario generation, the collected communication data and multi-modal sensory data are also deeply integrated and precisely aligned.
In summary, the developed platform integrates three high-precision software, i.e., AirSim, WaveFarer, and Wireless InSite, and further achieves in-depth integration and precise alignment of them. There are four steps involved in developing the simulation platform, including high-fidelity scenario construction (the first step), comprehensive scenario condition simulation (the second step), dynamic scenario generation (the third step), and data collection and export (the fourth step). Based on the developed platform, the SynthSoM dataset is constructed, which encompasses multiple data modalities under various air-ground multi-link cooperative scenarios with comprehensive conditions. For clarity, Table 2 depicts the key algorithms and parameters utilized in the construction of the SynthSoM dataset from the perspectives of four steps.
Software First Step Second Step Third Step Fourth Step
AirSim Rendering pipeline of UE4 Emit a multitude of layer rays AirSim plugin LISA algorithm Generate trajectories of dynamic intelligent agents based on SUMO and UE4 Export RGB images, depth maps, and LiDAR point clouds
WaveFarer Ray-tracing simulation for modeling radar paths Rain model Snow model Export mmWave radar data
Wireless InSite Ray-tracing simulation for modeling propagation paths Rain model Snow model Export communication channel data
Detailed parameter setting can be observed in https://github.com/ZiweiHuang96/SynthSoM.
Table 2: Key algorithms and parameters utilized in the construction of the SynthSoM dataset.
@@
The SynthSoM dataset is available at the Figshare repository[35]. To utilize the SynthSoM dataset for communication and multi-modal sensing tasks, we divide the SynthSoM dataset into five folders according to different scenarios. The scenario folder contains different conditions, and further contains different data modalities, i.e., RF sensing, non-RF sensing, and RF communications. The data directory hierarchy and content of the SynthSoM dataset are shown in Fig. 5. Due to the significantly massive data amount of the SynthSoM dataset, we utilize an efficient compression way and upload the SynthSoM dataset to the Figshare repository[35] in 7-zip compression format. Overall, the constructed SynthSoM dataset is fully open-sourced and has a size of 743.79 GB.
@@
Overall, the SynthSoM dataset contains 140K sets of channel matrices, 18K sets of path loss, 136K sets of mmWave radar waveforms with 38K radar point clouds, 145K RGB images, 290K depth maps, and 79K sets of LiDAR point clouds. The open-sourced SynthSoM dataset can provide consistent data for SoM-related algorithm cross comparison, model calibration, and baseline implementation.
@@
Figure 6: Visualization and statistical properties of RF channels under different conditions. (a) Path loss heatmap in urban crossroad scenario with sub-6 GHz frequency band. (b) Path loss heatmap in urban crossroad scenario with mmWave frequency band. (c) Propagation paths in mountain road scenario under SISO. (d) Propagation paths in mountain road scenario under MIMO. (e) Propagation paths in mountain road scenario under massive MIMO. (f) SCCFs in mountain road scenario with sub-6 GHz and mmWave frequency bands under MIMO and massive MIMO conditions. (g) Path loss heatmap in urban crossroad scenario on sunny days. (h) Path loss heatmap in urban crossroad scenario on rainy days. (i) Path loss heatmap in urban crossroad scenario on snowy days.